Connectivity Outage
Incident Report for SpecTronics UK
Resolved
The issue has now been resolved.
Posted May 11, 2020 - 18:10 BST
Monitoring
We have been notified by our upstream supplier that they have restored power to all impacted areas of their Data Centre as of 16:36.
Our NOC has completed full checks of our equipment located within the affected Data Centre and we can confirm that all affected circuits have now re-converged and traffic is passing as normal.

If you are experiencing any further issues, please contact our support team on 03333 444 114.
We sincerely apologise for the inconvenience caused.
Posted May 11, 2020 - 17:37 BST
Update
We are aware that a number of services are still being impacted due to the major power outage within the data centre.
We continue to liaise with our upstream supplier and understand the impact

A further update will be made available by 17:30 today
Posted May 11, 2020 - 16:37 BST
Identified
Our suppliers have confirmed that the majority of circuits have now returned to service, they are continuing to investigate but believe this to be related to a major power outage within a peering data centre.

A further update will be made available by 17:00 today
Posted May 11, 2020 - 15:44 BST
Investigating
Our upstream suppliers are currently experiencing a Major Incident affecting all connectivity products.

The incident started at 14:25 on 11/05/2020 and was confirmed as a priority one outage at 14:53. Customers with DSL and Ethernet services will be experiencing a total loss of service.

Our suppliers engineers are investigating and we will update by 16:00
Posted May 11, 2020 - 15:16 BST
This incident affected: Broadband/Connectivity Services (Broadband - ADSL, Broadband - FTTC, Leased Lines).